\begin{thebibliography}{10}
\providecommand{\url}[1]{#1}
\csname url@samestyle\endcsname
\providecommand{\newblock}{\relax}
\providecommand{\bibinfo}[2]{#2}
\providecommand{\BIBentrySTDinterwordspacing}{\spaceskip=0pt\relax}
\providecommand{\BIBentryALTinterwordstretchfactor}{4}
\providecommand{\BIBentryALTinterwordspacing}{\spaceskip=\fontdimen2\font plus
\BIBentryALTinterwordstretchfactor\fontdimen3\font minus
  \fontdimen4\font\relax}
\providecommand{\BIBforeignlanguage}[2]{{%
\expandafter\ifx\csname l@#1\endcsname\relax
\typeout{** WARNING: IEEEtran.bst: No hyphenation pattern has been}%
\typeout{** loaded for the language `#1'. Using the pattern for}%
\typeout{** the default language instead.}%
\else
\language=\csname l@#1\endcsname
\fi
#2}}
\providecommand{\BIBdecl}{\relax}
\BIBdecl

\bibitem{pontus:2012}
P.~Johnson, M.~Ekstedt, and I.~Jacobson, ``Where's the theory for software
  engineering?'' IEEE Software, vol.~29, no.~5, 2012, p.~96.

\bibitem{swebok:2014}
\BIBentryALTinterwordspacing
{IEEE Computer Society}, {Software Engineering Body of Knowledge (SWEBOK) -
  V3}, P.~Bourque and R.~Dupuis, Eds.\hskip 1em plus 0.5em minus 0.4em\relax
  EUA: Angela Burgess, 2014. [Online]. Available: \url{http://www.swebok.org/}
\BIBentrySTDinterwordspacing

\bibitem{bwMDD2010iu}
S.~L. Costa, V.~V. {Graciano Neto}, L.~F.~B. Loja, and J.~L. de~Oliveira, ``{ A
  Metamodel for Automatic Generation of Enterprise Information Systems},'' in
  {Proceedings of the First Brazilian Workshop on Model Driven Development
  (BW-MDD)}.\hskip 1em plus 0.5em minus 0.4em\relax Salvador, BA, Brazil: UFBA,
  2010, pp. 45--52.

\bibitem{bwMDD2010bpm}
L.~F.~B. Loja, V.~V. {Graciano Neto}, S.~L. Costa, and J.~L. de~Oliveira, ``{A
  Business Process Metamodel for Enterprise Information Systems Automatic
  Generation},'' in {Proceedings of the First Brazilian Workshop on Model
  Driven Development (BW-MDD)}.\hskip 1em plus 0.5em minus 0.4em\relax
  Salvador, BA, Brazil: UFBA, 2010, pp. 37--44.

\bibitem{earlyaspects:2011}
V.~V. {Graciano Neto} and J.~Lopes~de Oliveira, ``An early aspect for
  model-driven transformers engineering,'' in Proceedings of the 2011
  International Workshop on Early Aspects, ser. EA '11.\hskip 1em plus 0.5em
  minus 0.4em\relax Porto de Galinhas, Brazil: ACM, 2011, pp. 7--11.

\bibitem{bwMDD2011}
A.~H.~G. Oliveira, V.~V. {Graciano Neto}, S.~L. Costa, and J.~L. de~Oliveira,
  ``{Web Accessibility using Model-Driven Development},'' in {Proceedings of
  the II Brazilian Workshop on Model Driven Development (BW-MDD)}.\hskip 1em
  plus 0.5em minus 0.4em\relax Salvador, BA, Brazil: UFBA, 2011, pp. 18--25.

\bibitem{sbsi:2013}
V.~V. {Graciano Neto} and J.~L. de~Oliveira, ``Evolution of an application
  framework architecture for information systems with model-driven development
  (in portuguese),'' in SBSI 2013, vol.~1, Jo\~ao Pessoa, Brazil: UFPB, 2013, pp. 1--12.

\bibitem{sbsi:2014en}
S.~L. da~Costa, V.~V. {Graciano Neto}, B.~dos Reis~Calçado, and J.~L.
  de~Oliveira, ``{User Interface Stereotypes: A Model-Based Approach for
  Information Systems User Interfaces},'' in {Proceedings of the X Brazilian
  Symposium on Information Systems}, ser. SBSI '14.\hskip 1em plus 0.5em minus
  0.4em\relax Londrina, Brazil: UEL, 2014, pp. 1--12.

\bibitem{laweb:2014}
S.~L. da~Costa, V.~V. {Graciano Neto}, and J.~L. de~Oliveira, ``{A User
  Interface Stereotype to build Web Portals},'' in {Proceedings of the 9th
  Latin American Web Congress}, ser. LA-WEB '14.\hskip 1em plus 0.5em minus
  0.4em\relax Ouro Preto, Brazil: IEEE, 2014, pp. 1--9.

\bibitem{GracianoNeto2014b}
V.~V. Graciano~Neto, M.~Guessi, L.~B.~R. Oliveira, F.~Oquendo, and E.~Y.
  Nakagawa, ``Investigating the model-driven development for
  systems-of-systems,'' in Proceedings of the 2014 European Conference on
  Software Architecture Workshops, ser. ECSAW '14.\hskip 1em plus 0.5em minus
  0.4em\relax Vienna, Austria: ACM, 2014, pp. 22:1--22:8.

\bibitem{GracianoNeto:2015}
\BIBentryALTinterwordspacing
V.~V. Graciano~Neto, L.~Garc{\'e}s, M.~Guessi, L.~B.~R. de~Oliveira, and
  F.~Oquendo, ``On the equivalence between reference architectures and
  metamodels,'' in Proceedings of the 1st International Workshop on Exploring
  Component-based Techniques for Constructing Reference Architectures, ser.
  CobRA '15.\hskip 1em plus 0.5em minus 0.4em\relax New York, NY, USA: ACM,
  2015, pp. 21--24. [Online]. Available:
  \url{http://doi.acm.org/10.1145/2755567.2755572}
\BIBentrySTDinterwordspacing

\bibitem{GracianoNeto2016SRCPosBlind}
V.~V. {Graciano Neto}, ``Validating emergent behaviors in systems-of-systems
  through model transformations,'' in {ACM Student Research Competition at
  MODELS}, Saint Malo, France, October 2016, pp. 1--6.

\bibitem{GracianoNeto2017}
\BIBentryALTinterwordspacing
V.~V. Graciano~Neto, C.~E. Barros~Paes, L.~Garc{\'e}s, M.~Guessi, W.~Manzano,
  F.~Oquendo, and E.~Y. Nakagawa, ``Stimuli-sos: a model-based approach to
  derive stimuli generators for simulations of systems-of-systems software
  architectures,'' Journal of the Brazilian Computer Society, vol.~23, no.~1,
  Oct 2017, p.~13. [Online]. Available:
  \url{https://doi.org/10.1186/s13173-017-0062-y}
\BIBentrySTDinterwordspacing

\bibitem{GracianoNeto:HICSS2018}
V.~V. {Graciano Neto}, L.~Garc\'es, M.~Guessi, C.~Paes, W.~Manzano, F.~Oquendo,
  and E.~Nakagawa, ``{ASAS: An Approach to Support Simulation of Smart
  Systems},'' ser. The Hawaii International Conference on System Sciences
  (HICSS-51), Hawaii’s Big Island, USA, 2018, pp. 1--10, (Accepted for
  Publication).

\bibitem{DiRuscio:2014}
D.~D. Ruscio, R.~F. Paige, and A.~Pierantonio, ``Guest editorial to the special
  issue on success stories in model driven engineering,'' Science of Computer
  Programming, vol. 89, Part B, no.~0, 2014, pp. 69 -- 70.

\bibitem{Hutchinson:2014}
J.~Hutchinson, J.~Whittle, and M.~Rouncefield, ``Model-driven engineering
  practices in industry: Social, organizational and managerial factors that
  lead to success or failure,'' Science of Computer Programming, vol. 89, Part
  B, no.~0, 2014, pp. 144 -- 161, special issue on Success Stories in Model
  Driven Engineering.

\bibitem{buttner:2014}
F.~Buttner, U.~Bartels, L.~Hamann, O.~Hofrichter, M.~Kuhlmann, M.~Gogolla,
  L.~Rabe, F.~Steimke, Y.~Rabenstein, and A.~Stosiek, ``Model-driven
  standardization of public authority data interchange,'' Science of Computer
  Programming, vol. 89, Part B, no.~0, 2014, pp. 162 -- 175, special issue on
  Success Stories in Model Driven Engineering.

\bibitem{Cuadrado:2014}
J.~S. Cuadrado, J.~L.~C. Izquierdo, and J.~G. Molina, ``Applying model-driven
  engineering in small software enterprises,'' Science of Computer Programming,
  vol. 89, Part B, no.~0, 2014, pp. 176 -- 198, special issue on Success
  Stories in Model Driven Engineering.

\bibitem{sendall:2003}
S.~Sendall and W.~Kozaczynski, ``Model transformation: The heart and soul of
  model-driven software development,'' IEEE Software, vol.~20, no.~5, 2003, pp.
  42--45.

\bibitem{Hutchinson:2011}
J.~Hutchinson, J.~Whittle, M.~Rouncefield, and S.~Kristoffersen, ``Empirical
  assessment of mde in industry,'' in Proc. ICSE '11.\hskip 1em plus 0.5em
  minus 0.4em\relax Waikiki, Honolulu, HI, USA: ACM, 2011, pp. 471--480.

\bibitem{Hutchinson:2011b}
J.~Hutchinson, M.~Rouncefield, and J.~Whittle, ``Model-driven engineering
  practices in industry,'' in Proceedings of the 33rd International Conference
  on Software Engineering, ser. ICSE '11.\hskip 1em plus 0.5em minus
  0.4em\relax Waikiki, Honolulu, HI, USA: ACM, 2011, pp. 633--642.

\bibitem{amrani:2012}
M.~Amrani, L.~Lucio, G.~Selim, B.~Combemale, J.~Dingel, H.~Vangheluwe,
  Y.~Le~Traon, and J.~R. Cordy, ``A tridimensional approach for studying the
  formal verification of model transformations,'' in VOLT 2012, Montreal, QC,
  Canada, 2012, pp. 921--928.

\bibitem{bosems:2008}
S.~Bosems, ``Toward automated model transformations in the a-muse design
  methodology,'' in 9th Twente Student Conference on IT.\hskip 1em plus 0.5em
  minus 0.4em\relax University of Twente, 2008, pp. 1--7.

\bibitem{STVR:2014}
\BIBentryALTinterwordspacing
V.~Aranega, J.-M. Mottu, A.~Etien, T.~Degueule, B.~Baudry, and J.-L. Dekeyser,
  ``Towards an automation of the mutation analysis dedicated to model
  transformation,'' Software Testing, Verification and Reliability, 2014, pp.
  n/a--n/a. [Online]. Available: \url{http://dx.doi.org/10.1002/stvr.1532}
\BIBentrySTDinterwordspacing

\bibitem{cuadrado:2013}
J.~S. Cuadrado, E.~Guerra, and J.~de~Lara, ``The program is the model: Enabling
  transformations\@run.time,'' in Software Language Engineering, ser. Lecture
  Notes in Computer Science, K.~Czarnecki and G.~Hedin, Eds.\hskip 1em plus
  0.5em minus 0.4em\relax Springer Berlin Heidelberg, 2013, vol. 7745, pp.
  104--123.

\bibitem{conformsto}
J.~Canovas and J.~Molina, ``An architecture-driven modernization tool for
  calculating metrics,'' IEEE Software, vol.~27, no.~4, 2010, pp. 37--43.

\bibitem{Cicchetti:2008}
A.~Cicchetti, D.~D. Ruscio, R.~Eramo, and A.~Pierantonio, ``Automating
  co-evolution in model-driven engineering,'' in Proceedings of the 2008 12th
  International IEEE Enterprise Distributed Object Computing Conference.\hskip
  1em plus 0.5em minus 0.4em\relax Munich, Germany: IEEE Computer Society,
  2008, pp. 222--231.

\bibitem{weiss:2009}
G.~Weiss, G.~Pomberger, W.~Beer, G.~Buchgeher, B.~Dorninger, J.~Pichler,
  H.~Prahofer, R.~Ramler, F.~Stallinger, and R.~Weinreich, ``{Software
  Engineering and Processes and Tools},'' in Hagenberg Research, B.~Buchberger,
  M.~Affenzeller, A.~Ferscha, M.~Haller, T.~Jebelean, E.~P. Klement, P.~Paule,
  G.~Pomberger, W.~Schreiner, R.~Stubenrauch, R.~Wagner, G.~Wei{\ss}, and
  W.~Windsteiger, Eds.\hskip 1em plus 0.5em minus 0.4em\relax Springer Berlin
  Heidelberg, 2009, pp. 157--235.

\bibitem{smolander:2013}
K.~Smolander and T.~Paivarinta, ``Forming theories of practices for software
  engineering,'' in Software Engineering (GTSE), 2013 2nd SEMAT Workshop on a
  General Theory of, May 2013, pp. 27--34.

\bibitem{Cuadrado:2012}
J.~S\'{a}nchez-Cuadrado, J.~De~Lara, and E.~Guerra, ``Bottom-up meta-modelling:
  An interactive approach,'' in Proceedings of the 15th International
  Conference on Model Driven Engineering Languages and Systems.\hskip 1em plus
  0.5em minus 0.4em\relax Innsbruck, Austria: Springer-Verlag, 2012, pp. 3--19.

\bibitem{GuerraLWKKRSS13}
E.~Guerra, J.~de~Lara, M.~Wimmer, G.~Kappel, A.~Kusel, W.~Retschitzegger,
  J.~Sch{\"o}nb{\"o}ck, and W.~Schwinger, ``Automated verification of model
  transformations based on visual contracts,'' Autom. Softw. Eng., vol.~20,
  no.~1, 2013, pp. 5--46.

\bibitem{ambler:2003}
S.~Ambler, ``Agile model driven development is good enough,'' Software, IEEE,
  vol.~20, no.~5, Sept 2003, pp. 71--73.

\bibitem{basso:2010}
F.~P. Basso and R.~Pillat, ``An experience report on agile software development
  with mda (in portuguese: Um relato de experiência no desenvolvimento Ágil
  de software com a mda),'' in First Brazilian Workshop on Model Driven
  Development (I BW-MDD), Salvador, Brazil, 2010.

\bibitem{Oquendo:2006}
F.~Oquendo, ``{Pi-Method: A Model-driven Formal Method for Architecture-centric
  Software Engineering},'' SIGSOFT Softw. Eng. Notes, vol.~31, no.~3, May 2006,
  pp. 1--13.

\bibitem{oquendo07}
D.~Wu, J.~Chen, and F.~Oquendo, ``Formal model-driven engineering of
  distributed simulation systems based on architecture-centric domain-specific
  approach,'' in APSEC.\hskip 1em plus 0.5em minus 0.4em\relax Nagoya, Japan:
  ACM/IEEE, 2007, p. 568.

\bibitem{Ralph:2014}
\BIBentryALTinterwordspacing
P.~Ralph, I.~Exman, P.-W. Ng, P.~Johnson, M.~Goedicke, A.~T. Kocata, and K.~L.
  Yan, ``How to develop a general theory of software engineering: Report on the
  gtse 2014 workshop,'' SIGSOFT Softw. Eng. Notes, vol.~39, no.~6, Dec. 2014,
  pp. 23--25. [Online]. Available:
  \url{http://doi.acm.org/10.1145/2674632.2674647}
\BIBentrySTDinterwordspacing

\bibitem{jin:2001}
\BIBentryALTinterwordspacing
R.~Jin, W.~Chen, and T.~Simpson, ``\BIBforeignlanguage{English}{Comparative
  studies of metamodelling techniques under multiple modelling criteria},''
  \BIBforeignlanguage{English}{Structural and Multidisciplinary Optimization},
  vol.~23, no.~1, 2001, pp. 1--13. [Online]. Available:
  \url{http://dx.doi.org/10.1007/s00158-001-0160-4}
\BIBentrySTDinterwordspacing

\bibitem{essence:2001}
\BIBentryALTinterwordspacing
C.~Atkinson and T.~Kuhne, ``\BIBforeignlanguage{English}{The essence of
  multilevel metamodeling},'' in \BIBforeignlanguage{English}{UML 2001 - The
  Unified Modeling Language. Modeling Languages, Concepts, and Tools}, ser.
  Lecture Notes in Computer Science, M.~Gogolla and C.~Kobryn, Eds.\hskip 1em
  plus 0.5em minus 0.4em\relax Springer Berlin Heidelberg, 2001, vol. 2185, pp.
  19--33. [Online]. Available: \url{http://dx.doi.org/10.1007/3-540-45441-1_3}
\BIBentrySTDinterwordspacing

\bibitem{Hug2009}
\BIBentryALTinterwordspacing
C.~Hug, A.~Front, D.~Rieu, and B.~Henderson-Sellers, ``A method to build
  information systems engineering process metamodels,'' Journal of Systems and
  Software, vol.~82, no.~10, 2009, pp. 1730 -- 1742, sI: \{YAU\}. [Online].
  Available:
  \url{http://www.sciencedirect.com/science/article/pii/S0164121209001174}
\BIBentrySTDinterwordspacing

\bibitem{deLara:2010}
\BIBentryALTinterwordspacing
J.~de~Lara and E.~Guerra, ``Deep meta-modelling with metadepth,'' in
  Proceedings of the 48th International Conference on Objects, Models,
  Components, Patterns, ser. TOOLS'10.\hskip 1em plus 0.5em minus 0.4em\relax
  Berlin, Heidelberg: Springer-Verlag, 2010, pp. 1--20. [Online]. Available:
  \url{http://dl.acm.org/citation.cfm?id=1894386.1894387}
\BIBentrySTDinterwordspacing

\bibitem{Li:2010}
\BIBentryALTinterwordspacing
Y.~F. Li, S.~H. Ng, M.~Xie, and T.~N. Goh, ``A systematic comparison of
  metamodeling techniques for simulation optimization in decision support
  systems,'' Appl. Soft Comput., vol.~10, no.~4, Sep. 2010, pp. 1257--1273.
  [Online]. Available: \url{http://dx.doi.org/10.1016/j.asoc.2009.11.034}
\BIBentrySTDinterwordspacing

\bibitem{Cicchetti466155}
A.~Cicchetti, D.~Di~Ruscio, D.~S. Kolovos, and A.~Pierantonio, ``A test-driven
  approach for metamodel development,'' in Emerging Technologies for the
  Evolution and Maintenance of Software Models.\hskip 1em plus 0.5em minus
  0.4em\relax Business science reference, 2011, pp. 319--342.

\bibitem{Bak:2013}
\BIBentryALTinterwordspacing
K.~B\k{a}k, D.~Zayan, K.~Czarnecki, M.~Antkiewicz, Z.~Diskin, A.~W\k{a}sowski,
  and D.~Rayside, ``Example-driven modeling: Model = abstractions + examples,''
  in Proceedings of the 2013 International Conference on Software Engineering,
  ser. ICSE '13.\hskip 1em plus 0.5em minus 0.4em\relax Piscataway, NJ, USA:
  IEEE Press, 2013, pp. 1273--1276. [Online]. Available:
  \url{http://dl.acm.org/citation.cfm?id=2486788.2486982}
\BIBentrySTDinterwordspacing

\bibitem{BergmayrW13}
\BIBentryALTinterwordspacing
A.~Bergmayr and M.~Wimmer, ``Generating metamodels from grammars by chaining
  translational and by-example techniques,'' in Proceedings of the First
  International Workshop on Model-driven Engineering By Example co-located with
  {ACM/IEEE} 16th International Conference on Model Driven Engineering
  Languages and Systems (MoDELS 2013), Miami, Florida, USA, September 29,
  2013., 2013, pp. 22--31. [Online]. Available:
  \url{http://ceur-ws.org/Vol-1104/3.pdf}
\BIBentrySTDinterwordspacing

\bibitem{ArianNik2014}
\BIBentryALTinterwordspacing
M.~A. Nik, K.~Fayazbakhsh, D.~Pasini, and L.~Lessard, ``A comparative study of
  metamodeling methods for the design optimization of variable stiffness
  composites,'' Composite Structures, vol. 107, no.~0, 2014, pp. 494 -- 501.
  [Online]. Available:
  \url{http://www.sciencedirect.com/science/article/pii/S0263822313004200}
\BIBentrySTDinterwordspacing

\bibitem{ClarkSW15}
\BIBentryALTinterwordspacing
T.~Clark, P.~Sammut, and J.~S. Willans, ``Applied metamodelling: {A} foundation
  for language driven development (third edition),'' CoRR, vol. abs/1505.00149,
  2015. [Online]. Available: \url{http://arxiv.org/abs/1505.00149}
\BIBentrySTDinterwordspacing

\bibitem{Sen2015}
\BIBentryALTinterwordspacing
O.~Sen, S.~Davis, G.~Jacobs, and H.~Udaykumar, ``Evaluation of convergence
  behavior of metamodeling techniques for bridging scales in multi-scale
  multimaterial simulation,'' Journal of Computational Physics, vol. 294,
  no.~0, 2015, pp. 585 -- 604. [Online]. Available:
  \url{http://www.sciencedirect.com/science/article/pii/S0021999115001989}
\BIBentrySTDinterwordspacing

\bibitem{Piho2014457}
\BIBentryALTinterwordspacing
G.~Piho, J.~Tepandi, D.~Thompson, T.~Tammer, M.~Parman, and V.~Puusep,
  ``Archetypes based meta-modeling towards evolutionary, dependable and
  interoperable healthcare information systems,'' Procedia Computer Science,
  vol.~37, no.~0, 2014, pp. 457 -- 464, the 5th International Conference on
  Emerging Ubiquitous Systems and Pervasive Networks (EUSPN-2014)/ The 4th
  International Conference on Current and Future Trends of Information and
  Communication Technologies in Healthcare (ICTH 2014)/ Affiliated Workshops.
  [Online]. Available:
  \url{http://www.sciencedirect.com/science/article/pii/S1877050914010345}
\BIBentrySTDinterwordspacing

\bibitem{lncs:2014}
\BIBentryALTinterwordspacing
G.~M. Selim, L.~Lucio, J.~R. Cordy, J.~Dingel, and B.~J. Oakes,
  ``\BIBforeignlanguage{English}{Specification and verification of graph-based
  model transformation properties},'' in \BIBforeignlanguage{English}{Graph
  Transformation}, ser. Lecture Notes in Computer Science, H.~Giese and
  B.~Konig, Eds.\hskip 1em plus 0.5em minus 0.4em\relax Springer International
  Publishing, 2014, vol. 8571, pp. 113--129. [Online]. Available:
  \url{http://dx.doi.org/10.1007/978-3-319-09108-2_8}
\BIBentrySTDinterwordspacing

\bibitem{LaraG05}
J.~de~Lara and E.~Guerra, ``Formal support for model driven development with
  graph transformation techniques,'' in DSDM - Agile 2005.\hskip 1em plus 0.5em
  minus 0.4em\relax Denver, Colorado: CEUR, 2005.

\bibitem{gramot2005_tax}
T.~Mens, P.~V. Gorp, G.~Karsai, and D.~Varr{\'o}, ``Applying a model
  transformation taxonomy to graph transformation technology,'' in GraMot 2005,
  International Workshop on Graph and Model Transformations, ser. ENTCS,
  G.~Karsai and G.~Taentzer, Eds., vol. 152, Elsevier.\hskip 1em plus 0.5em
  minus 0.4em\relax Tallinn, Estonia: Elsevier, 2006, pp. 143--159.

\bibitem{Schuerr2008}
A.~Schurr, M.~Nagl, and A.~Zundorf, Eds., Applications of Graph Transformations
  with Industrial Relevance.\hskip 1em plus 0.5em minus 0.4em\relax Berlin:
  Springer-Verlag, 2008.

\bibitem{Greenyer2012}
J.~Greenyer and J.~Rieke, ``Applying advanced tgg concepts for a complex
  transformation of sequence diagram specifications to timed game automata,''
  in Proceedings of the 4th International Conference on Applications of Graph
  Transformations with Industrial Relevance (AGTIVE 2011), ser.
  AGTIVE'11.\hskip 1em plus 0.5em minus 0.4em\relax Budapest, Hungary:
  Springer-Verlag, 2012, pp. 222--237.

\bibitem{Czarnecki:2003}
K.~Czarnecki and S.~Helsen, ``Classification of model transformation
  approaches,'' in Proc. of OOPSLA’03 Workshop on Generative Techniques in
  the Context of Model-Driven Architecture, Anaheim, California, USA, 2003.

\bibitem{Cho:2011}
\BIBentryALTinterwordspacing
H.~Cho and J.~Gray, ``Design patterns for metamodels,'' in Proceedings of the
  Compilation of the Co-located Workshops on DSM'11, TMC'11, AGERE! 2011,
  AOOPES'11, NEAT'11, \& VMIL'11, ser. SPLASH '11 Workshops.\hskip 1em plus
  0.5em minus 0.4em\relax New York, NY, USA: ACM, 2011, pp. 25--32. [Online].
  Available: \url{http://doi.acm.org/10.1145/2095050.2095056}
\BIBentrySTDinterwordspacing

\bibitem{ortiz:2014}
\BIBentryALTinterwordspacing
F.~J. Ortiz, D.~Alonso, F.~Rosique, F.~Sanchez-Ledesma, and J.~Pastor,
  ``\BIBforeignlanguage{English}{A component-based meta-model and framework in
  the model driven toolchain c-forge},'' in
  \BIBforeignlanguage{English}{Simulation, Modeling, and Programming for
  Autonomous Robots}, ser. Lecture Notes in Computer Science, D.~Brugali,
  J.~Broenink, T.~Kroeger, and B.~MacDonald, Eds.\hskip 1em plus 0.5em minus
  0.4em\relax Springer International Publishing, 2014, vol. 8810, pp. 340--351.
  [Online]. Available: \url{http://dx.doi.org/10.1007/978-3-319-11900-7_29}
\BIBentrySTDinterwordspacing

\bibitem{Ghani:2014:HRE}
I.~Ghani, Handbook of Research on Emerging Advancements and Technologies in
  Software Engineering, 1st~ed.\hskip 1em plus 0.5em minus 0.4em\relax Hershey,
  PA, USA: IGI Global, 2014.

\end{thebibliography}

\documentclass[a4paper]{article}

\usepackage[english]{babel}
\usepackage[utf8x]{inputenc}
\usepackage[T1]{fontenc}

\usepackage[a4paper,top=3cm,bottom=2cm,left=3cm,right=3cm,marginparwidth=1.75cm]{geometry}

\usepackage{amsmath}
\usepackage{graphicx}
\usepackage[colorinlistoftodos]{todonotes}
\usepackage[colorlinks=true, allcolors=blue]{hyperref}

\title{Towards Methods for Model-Based Software Development}
\author{Valdemar Vicente Graciano Neto and
Luiz Fernando Batista Loja\thanks{Instituto de Inform\'atica (INF/UFG), Universidade Federal de Goi\'as, Goi\^ania, Brazil; ICMC, University of S\~ao Paulo (USP), S\~ao Carlos, Brazil;Instituto Federal de Goi\'as (IFG), Luzi\^ania, Brazil. Email: valdemarneto@inf.ufg.br, luizloja@gmail.com}
}

\begin{document}
\maketitle

\begin{abstract}
Software engineering is a young discipline. Despite efforts in recent decades, some elements still require further development, research, and systematization. One of these elements are methods for software engineering. They consist of a set of well-defined activities used in a software development to guide how the work should be performed to achieve the expected results. However, there is a lack of systematic knowledge that effectively guide how work should be done in a variety of areas. Despite the knowledge available in Software Engineering Body of Knowledge (SWEBOK), only the classical methods are comprised there, while emerging topics such as Agent-Oriented Software Engineering (AOSE), Aspect-Oriented Software Development (AOSD), and Model-Based Software Development (MBSE) are only briefly mentioned. We claim that it is also necessary to investigate methods for those emerging software engineering subtracks. In this direction, this paper presents a a speculative and preliminary emerging results on the establishment of methods for MBSE. We report some insights on the conception of methods for MBSE.
\end{abstract}

\section{Introduction}

Software Engineering is a young engineering. Despite the success achieved in the last decades, the abstract nature of software has been challenging for predictions of cost, time, quality. There is also a lack of systematization on procedures, processes, and methods to guide a successful software development. Software Engineering Body of Knowledge (SWEBOK) is one among other attempts to systematize the Software Engineering's knowledge \cite{pontus:2012, swebok:2014}. Most of the Software Engineering subareas are covered there. However, emerging paradigms are not covered, yet. Feature-Driven Development (FDD) and Model-Based Software Development (MBSE) are only briefly cited, while Agent-Oriented Software Engineering (AOSE) is not mentioned \cite{swebok:2014}. Software  engineering  methods  provide  an  organized and systematic approach to developing software. There are numerous methods  to choose, and it is important  for a software engineer to choose an appropriate one or a set of them for the development at hand. Nevertheless, such choice has a dramatic effect on the success of the software project \cite{swebok:2014}. 

Despite the importance of such topic, there is a lack of details on how they should be structured, which recurrent activities comprise them, how they are related to the software development processes, and how they could be associated to produce a better result. Thus, in those situations, software engineers come back again to the pragmatism and experience-based decisions, since there is no a theoretical basis to ground their decisions. SEMAT (Software Engineering Method and Theory)\footnote{\url{http://semat.org/}} initiative is a remarkable instance of means to systematize methods and knowledge in software engineering. However, for recent areas such as MBSE, there is still a lack of systematization on methods. 

We have invested efforts on MBSE in the last years  \cite{bwMDD2010iu,bwMDD2010bpm,earlyaspects:2011,bwMDD2011, sbsi:2013, sbsi:2014en, laweb:2014, GracianoNeto2014b, GracianoNeto:2015, GracianoNeto2016SRCPosBlind, GracianoNeto2017, GracianoNeto:HICSS2018}. Despite MBSE successful cases \cite{DiRuscio:2014, Hutchinson:2014, buttner:2014, Cuadrado:2014}, there is still a lack of consensual set of techniques to develop a metamodel, for example, or to engineer a bidirectional transformation. Indeed, in software engineering, one of the most hotly debated questions concerns the choice of method \cite{pontus:2012}. And, when you reduce this scope for MBSE, we claim that the same occurs. 

This paper presents preliminary insights on methods for MBSE. We present a brief discussion on the basis which could be followed to edify a Methods Theory for MBSE. Aiming at suitably present the content, this paper is structured as follows: Section II presents some foundations to foster and feed the discussion, Section III presents our initial perceptions on the topic, and Section IV presents a brief discussion, while Section V brings final considerations and discusses future work perspectives.

\section{Background}


Essentially, a software development process produces models that are successively transformed along the process until reaching software code. In this scenario, Model-Based Software Development (MBSE) emerged as a solution where models are the first-class citizen \cite{sendall:2003, Hutchinson:2011, Hutchinson:2011b, amrani:2012}. Model transformations are performed in an automated way, through the use of specific software tools \cite{bosems:2008, STVR:2014}. Now, the program is the model \cite{cuadrado:2013}
Adopting MBSE requires artifacts and tools. In MBSE, remarkable artifacts involve models, metamodels, model transformers, and model transformations. A \textbf{model} is, essentially, an abstraction of reality represented in a textual or visual language (usually, a Domain-Specific Language (DSL)). \textbf{Metamodels} are special models that guides and restricts how to construct other models. Models necessarily need to conform to their respective metamodels \cite{conformsto, Cicchetti:2008}. \textbf{Model transformers} are considered model compilers which receive models and their respective metamodels as input and, through the use of \textbf{model transformations}, they transform source models in target models. 	MBSE can be considered one of the innovative software development processes recently created. As such, MBSE is also based on processes, methods, activities, and technologies which support them.


In parallel, processes, methods, techniques, and activities are part of Software Engineering. A process is a framework which involves methods, techniques, activities, and tools \cite{swebok:2014}. A process can have many methods associated (i.e., a process can have Object-Oriented method and Agile methods applied in parallel). A method can be a part of zero or many processes (i.e., formal methods are used in some processes, but many software development processes do not use it). One method has at least one inherent technique (to exist), or many techniques related to it (i.e., classes identification via grammar classes for OO-method, stand-up meetings for agile methods, formal specification for formal methods). And one technique can be part of many processes, as domain engineering, that analyses commonalities and variabilities, and can be a part of feature-driven development, or product-line software development.

Methods should be enough to predict in the each stage of software developments the quality of the end product and respective techniques which should be adopted to achieve success and to avoid repetitive labor and error \cite{pontus:2012}. However, this is not true yet for software engineering, specially for MBSE. In SWEBOK, methods theory is structured in four well-delimited categories:  Heuristic Methods (Structured paradigm, data modeling paradigm, and object-oriented paradigm), Formal Methods (program specification and validation, and program refinement and derivation), Prototyping Methods (styles, target, and evaluation techniques), and Agile Methods (RAD, XP, Scrum, and FDD) \cite{swebok:2014}. Significant progress has been made over the past decades in establishing methods and techniques that support the demands on software development processes. However, there is still a need for improving those methods and techniques \cite{weiss:2009}, besides proposing new ones for emerging paradigms such as MBSE.


%






\section{Methods for MBSE}

For have observed pragmatic issues and acquired experience on MBSE in the last years. With such experience, we tried to gather part of this to structure and  potentially generalize our knowledge \cite{smolander:2013}. A first step to characterize methods is exposing the inherent techniques and activities that recurrently compose them. When considering techniques, they are usually embodied in a method, and they are supported by tools \cite{swebok:2014}. From the best of our knowledge, there is a lack of studies that characterize MBSE techniques, and group them in methods. Hence, we struggle to elicit a set of techniques which characterizes a \textit{Model-Driven Method (MDM)} for each potential stage of a software development process. Then, we can catalog a set of techniques to characterize a MDM for both flows: to construct MBSE tools, and another to use MBSE tools.

The main activity of MBSE is \textit{modeling}. Consequently, every technique related to modeling activity could be added to a MDM. Another important activity in MBSE is \textit{metamodeling}. Similarly, every technique related to metamodeling activity could be added to a MBSE method. Then, we could list at least these main activities in MBSE: modeling, metamodeling \cite{Cuadrado:2012}, checking consistency between model and metamodel \cite{GuerraLWKKRSS13}, transforming models to models, and transforming models to code. 

Tables \ref{tab:process1} and \ref{tab:process2} summarize what could be an MDM: Table \ref{tab:process1} shows the MBSE tools engineering flow, its activities, and each technique that characterize the MDM for that activity. Table \ref{tab:process2} shows the techniques that characterizes MDM use in a classic software development process for each activity of a development process.

Modeling in this context is the activity of \textit{instantiating} a metamodel to obtain a concrete model. And metamodeling is the activity of performing a domain engineering to elicit and document all the concepts that characterizes an specific domain that should be included in a conceptual model for that domain, in such a way that concrete models could be \textit{instantiated} from them, keeping the conformance relation.

\begin{table}

	\centering
		\begin{tabular}{|p{5cm}|p{6cm}|}
		  \hline \textbf{Activity} & \textbf{Techniques} \\\hline
			 Metamodel Specification & Domain Engineering  \\ \hline
			 Metamodel Design & Separation of Concerns in specific models, Concepts representation   \\ \hline
			 Metamodel Implementation & Converting in a textual language (tangled with a transformer implementation, or modularized in a XML, or other structure)  \\\hline 
			 Model Transformation Engineering & Linking one or more source metamodel concepts to one or more target metamodel concepts. \\\hline
			\end{tabular}
	\caption{MBSE engineering flow and its techniques.}
	\label{tab:process1}
\end{table}

\begin{table}
	\centering
		\begin{tabular}{|p{5cm}|p{6cm}|}
		  \hline \textbf{Activity} & \textbf{Techniques} \\\hline
			 Requirements Engineering & Modeling (instantiating a metamodel), checking consistency between model (with tool support or not)  \\ \hline
			 Design & Transform models to models   \\ \hline
			 Codification & Transform models to code  \\\hline 
			 Testing & Using models to extract test cases and test them \\\hline
			\end{tabular}
	\caption{Application Engineering and MDM techniques}
	\label{tab:process2}
\end{table}


It is important to highlight that tables present only techniques that could be considered inherent to MDM application. Analogously to mentioned before, other techniques from other methods as Object-Oriented, Agile, for Formal could be used in association with MDM application.

\section{Discussion}

MBSE has not enough maturity about a taxonomy of methods and techniques, and efforts has been performed to establish MBSE as a complete software development process, with a set of methods and their respective set of techniques.  SWEBOK \cite{swebok:2014} defines a method as an organized and systematic approach for developing software for a target computer. According to SWEBOK, methods can be classified as Heuristic Methods (Procedural, Object-Oriented, and Data Modeling), Formal Methods, Prototyping Methods, and Agile Methods. From the best of our knowledge, there is no study in the literature that characterizes MBSE as a method in the strict meaning of the word in Software Engineering under SWEBOK perspective. 

Considering that classification, MBSE can be considered itself as new method that is not mentioned yet because did not reach enough maturity (in a set that includes Agent-Oriented Software Development, Aspect-Oriented Software Development, Feature-Oriented Software Development, and any another new technology that originates new software development paradigms). On the other hand, there some occurrences of associations between these main method areas and MBSE, as Agile-MBSE \cite{ambler:2003, basso:2010}, and Formal MBSE \cite{Oquendo:2006, oquendo07}.  Additionally, there is no study that consider MBSE as a software engineering method. MBSE is closer to the context of software development process \cite{sbsi:2013}. It involves processes, methods, and techniques, as other software development process. However, we understand that most of the software development life cycle has a method associated: Agile and agile method, and prototyping and prototyping method. In fact, there is still a need for a consensual vocabulary to describe issues such as paradigm, method, process, technique, approach, methodology, since those words often are used interchangeably.

Mature  academic  disciplines  typically  emphasize  the  importance  of their central theories \cite{Ralph:2014}. Software engineering (SE), in contrast, has no  widely  accepted  central  or  general  theories  \cite{pontus:2012}. SE  would  benefit  from  a  general  theory  in  numerous  ways \cite{Ralph:2014}: 1)  general  theories facilitate  developing  a  cumulative,  communicable  body  of  knowledge by preserving existing results and implicitly coordinating future studies; 2)  general  theories  improve  resistance  to  fads  by  stabilizing  the discipline; 3) general theory helps root out dubious and pseudoscientific concepts. 

Some research paths can be visualized regarding a consolidation of a Method Theory for MBSE:

\begin{itemize}
	\item \textbf{A consolidation of standard techniques for metamodeling}. Metamodeling is a core recurrent activity in MBSE. However, there is no consensus on the best techniques to perform it since several authors have proposed their own guidelines to construct metamodels \cite{jin:2001, essence:2001, Hug2009, deLara:2010, Li:2010, Cicchetti466155, Bak:2013, BergmayrW13, ArianNik2014, ClarkSW15, Sen2015, Piho2014457}. An investigation on this topic can foster methods consolidation for MBSE since its methods intensively depends on such activity;
	\item \textbf{A unified recommendation on how to transform models}. Some paradigms also are known for model transformations \cite{lncs:2014}. 	Model-to-model (M2M) transformations usually use graph patterns \cite{LaraG05, gramot2005_tax, Schuerr2008, Greenyer2012}. 	Model-to-text (or Model-to-code (M2C))  category is distinguished  between  visitor-based  and  template-based approaches \cite{Czarnecki:2003}. In former, a very basic code generation approach consists in providing some visitor mechanism to traverse the  internal  representation  of  a  model  and  write  code  to  a  text  stream to generate code. In latter, a template of a program in the target technology is filled with code transformed from a source model. Further research is necessary to establish a unifying approach for model transformation, regardless of the approach used;
\end{itemize}

Other activities need to be better specified regarding MBSE, such as the application of design patterns for metamodeling \cite{Cho:2011}, and the association with Component-Based Software Engineering \cite{ortiz:2014}. However, we believe that consolidating metamodeling and transformation techniques could properly benefit several other associations of MBSE with other initiatives, since metamodeling and model transformation are the heart and soul of MBSE \cite{sendall:2003}. In fact, several methods for MBSE have been proposed in the literature \cite{Ghani:2014:HRE}. MDA itself can be considered an specific MBSE method. However, MDA for examples lacks from recommendations on how to engineer a transformer in such a way that its architecture supports changes in the transformation rules without impacts for the transformer's architecture \cite{earlyaspects:2011}. Other initiatives do exist. The diversity of such distinct initiatives is also a claiming for a unifying methods theory for MBSE.  

\section{Final Remarks}\label{sec:MBSEfinal}

In this paper, we reported preliminary insights on the investigation of methods for Model-Based Software Development (MBSE). We claim that MBSE does not hold a method theory and that, with the realization of this for MBSE, well-grounded results could be generalized to the dawn of a general theory of methods for software engineering. That theory would integrate a future \textit{everything theory for software engineering}. This is only an embryonic reflexion to instigate the software engineering community on the emergency of this topic. Further research needs to be performed to suitably ground a theory for methods in MBSE, and for software engineering as a whole.

\bibliographystyle{alpha}

\end{document}